OpenStudy (jhulian):

solve this 2x^2 - 1 = 0

OpenStudy (jhulian):

@kweenlexis

OpenStudy (anonymous):

add 1

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

take the square root, don't forget the \(\pm\)

OpenStudy (jhulian):

thanks...but does it look like difference of two squares?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

un less you want to write it as \[(\sqrt{2}x+1)(\sqrt{2}x-1)=0\]